00Bastian Gruberx

 00  x
 00  x

00Professional x 00LinkedIn: linkedin.com/in/bastiangruberx 00Book: rustwebdevelopment.com x 00E-Mail: gruberbastian@me.com x 00Website: bastiangruber.ca x 00Code x 00GitHub: github.com/gruberb x 00Git: code.novanexus.ca/gruberbx

SUMMARY

Staff Software Engineer at Mozilla, shipping Rust, C++, and Python to hundreds of millions of Firefox users. Published author (“Rust Web Development” – Manning). Founder of Rust&Tell Berlin, the largest Rust meetup globally.

SKILLS

Languages: Rust, C++, Python, TypeScript, Go Backend & Systems: Rust (Actix, Warp, Axum), Python (FastAPI), Node.js, async Rust, FFI/UniFFI

Patterns: REST, GraphQL, gRPC, distributed systems (P2P), microservices Infrastructure: AWS, Google Cloud, Docker, Kubernetes, CI/CD, Grafana AI Tooling: Claude Code, GitHub Copilot

PROFESSIONAL EXPERIENCE

Staff Software Engineer (promoted from Senior) July 2024 - Present

Mozilla, Remote

Implemented Oblivious HTTP (OHTTP) from scratch in application-services (shared Rust component cross-compiled to Swift/Kotlin via UniFFI) plus C++ and Rust for Firefox Desktop – users are now unidentifiable when making privacy-sensitive requests across all platforms

Consolidated four separate favicon implementations into a single Image Service – extended an existing Python/Airflow pipeline rather than building a new microservice, now serving 4 teams from one manifest behind a GCS-backed API

Maintained Firefox Push notification service (50k req/s, sub-200ms latency): reduced memory footprint, added circuit breakers, built Grafana dashboards

Contributor and code reviewer for UniFFI, Mozilla’s multi-language bindings generator (Rust to Kotlin, Swift, Python, JavaScript)

Led “Rusty Fridays” internal education sessions, mentoring engineers across teams through Rust adoption

Senior Systems Engineer, Team Lead 2023 - 2024

Toposware, Remote Core team of 4 building a peer-to-peer distributed network protocol from scratch in Rust.

Achieved 10x throughput per node by profiling the actual bottleneck – optimized batch processing and serialization paths rather than redesigning the consensus layer

Led benchmark initiative that revealed nodes were compute-bound on crypto operations, not network-limited – redirected optimization effort accordingly

Built an EVM bridge connecting external Ethereum-compatible chains to the core network

Senior Software Engineer 2022 - 2023

Centrifuge, Remote Protocol team building DeFi applications on Substrate/Polkadot in Rust.

Implemented on-chain fund pool mutations – core feature for the lending protocol enabling live capital allocation adjustments

Replaced manual documentation with CI-driven auto-generation from Rust source – eliminated persistent drift between code and public docs

Author, “Rust Web Development” 2020 - 2022

Manning Publications Published end-to-end guide covering async Rust, web frameworks, and production deployment.

Sr. Solutions Architect (promoted from SA) 2020 - 2022

Twilio Inc., Remote

Technical lead for Fortune 500 engagements: debugged latency issues, redesigned messaging flows, advised on architectures matching client team size and resources

Created reusable engagement playbooks adopted across the EMEA Professional Services team Systems Engineer 2019 - 2020

Kraken Digital Currency Exchange, Remote

Migrated core banking services and API gateway from PHP to Rust, achieving 100x throughput improvement – enabling order-of-magnitude traffic growth without scaling hardware

Worked directly with the CTO on privacy-focused architectural decisions for the trading platform

Earlier Career 2014 - 2019

Sr. Software Engineer across enterprise clients (OSRAM, Kxrber Digital, GIATA, Avenga) and startups (Dalia Research, blogfoster, smartB) in Berlin.

Founded “Rust & Tell Berlin” meetup – grew it into the largest Rust meetup globally, still running years later

Designed IoT authentication infrastructure at OSRAM for thousands of connected devices (Node.js, Go, Lua)

First engineering hire at Kxrber Digital: built real-time sensor data pipelines on AWS

Decomposed GIATA’s monolithic ERP into 6 Docker-based microservices, improving system scalability

EDUCATION

Bachelor of Science, Computer Science TH Ingolstadt – Scholarship from Allianz Deutschland AG (fully funded)